Galactosemia is a recessive human disease that is treatable by restricting lactose and glucose in the diet. Susan Smithers and her husband are both heterozygous for the galactosemia gene.
Answer #1

C). Given that both the parents are heterozygous. Cross between them will result in the offspring with the following genotypes.

Aa* Aa -----> A_ (3/4, normal), aa (1/4, affected)

The probability of having all 4 four children with normal phenotypes is, = (3/4)^4 = 81/256

d). The probability of having at least one affected children is = 1 - probability of having no affected children = 1-81/256 = 175/256.